OpenStudy (ny,ny):

pleaseeee helps solve and check: b-4/b-2 = b-2/b+2 + 1/b-2

OpenStudy (ny,ny):

\[\frac{ b-4 }{ b-2 } =\frac{ b-2 }{ b+2 } + \frac{ 1 }{ b-2 }\]

OpenStudy (ny,ny):

@Compassionate

OpenStudy (compassionate):

Are you trying to solve for b?

OpenStudy (ny,ny):

yeah i think so

OpenStudy (compassionate):

First find the greatest common factor. Do you know what that is?

OpenStudy (ny,ny):

umm would that (b-2)(b+2)?

OpenStudy (ny,ny):

would that be*

OpenStudy (compassionate):

Correct. Now, multiply that by each numerator, but exclude the term if it is already in the numerator.

OpenStudy (ny,ny):

so instead of (b-2)(b+2) times (b-2) do just (b+2) times (b-2)?

OpenStudy (compassionate):

\(\frac{ b-4 }{ b-2 } =\frac{ b-2 }{ b+2 } + \frac{ 1 }{ b-2 }\) \(\frac{ b-4(b+2)}{ b-2 } =\frac{ b-2 }{ b+2 } + \frac{ 1 }{ b-2 }\) Notice that we multiplied it by b + 2, and not b - 2, this is because b - 2 is in its denominator.

OpenStudy (compassionate):

You will do this for every term.

OpenStudy (ny,ny):

b^2 + 6b +8 = b^2 - 4b + 4 + b +2

OpenStudy (compassionate):

Now that you have it eliminated, solve like a normal equation

OpenStudy (ny,ny):

b^2 + 6b + 8 = b^2 -3b + 6

OpenStudy (compassionate):

Continue.

OpenStudy (ny,ny):

subtract b^2 both sides?

OpenStudy (compassionate):

Nope. Subtract 6.

OpenStudy (ny,ny):

b^2 = b^2 + 9b + 2

OpenStudy (compassionate):

b^2 + 6b + 8 = b^2 -3b + 6 b^2 + 6b = b^2 - 3b + 6 b^2 = b^2 - 9b + 6

OpenStudy (ny,ny):

whered the 8 go?

OpenStudy (compassionate):

b^2 + 6b + 8 = b^2 -3b + 6 I messed up. We should've subtracted 6, which wouldn've maken it 2.

OpenStudy (ny,ny):

okay which gives us b^2 + 6b + 2 = b^2-3b

OpenStudy (ny,ny):

@Compassionate

OpenStudy (compassionate):

Now subtract the 2 Add the 3b Divide..

OpenStudy (ny,ny):

b^2 + 9b = b^2 - 2 divide the nine? b^2 + b = b^2/9 - 2/9 ? @Compassionate